Verify that the given family of functions solves the differential equation.
dy/dt = (1-2t) y² , y = 1 / (C - t + t² )

dy/dt = y² sin(t) , y = 1 / (C +cos(t))
1
Expert's answer
2020-07-02T19:30:23-0400

"\dfrac{dy}{dt}=(1-2t)y^2\newline
\dfrac{dy}{y^2}=(1-2t)*dt\newline
\dfrac{-1}{y}=t-t^2-C\newline
y=\dfrac{1}{(C-t+t^2)}\newline
\dfrac{dy}{dt}=y^2*\sin{(t)}\newline
\dfrac{dy}{y^2}=\sin{(t)}*dt\newline
\dfrac{-1}{y}=-\cos{(t)}-C\newline
y=\dfrac{1}{C+\cos{(t)}}"